[LA] Check that {1,(x+1),(x+1)^2} is a basis of the vector space of polynomial over R of degree at most 2. Find the coordinate of 3+x+2x^2 with respect to the basis.


1
Expert's answer
2021-12-22T16:02:41-0500

1. Consider the standard basis "B=\{1, x, x^2\}" on "P_2." Using this basis, we can write the elements using coordinate vectors as

"[1]_B=\begin{bmatrix}
 1\\
 0 \\
0
\end{bmatrix},[x+1]_B=\begin{bmatrix}
 1\\
 1 \\
0
\end{bmatrix},[(x+1)^2]_B=\begin{bmatrix}
 1\\
 2 \\
1
\end{bmatrix}"

"(x+1)^2=1+2x+x^2"

Because "\dim P_2=3," this set is a basis if and only if these three vectors are linearly independent. To verify this, consider the matrix


"\begin{pmatrix}
 1 & 1 & 1 \\
 0 & 1 & 2 \\
0 & 0 & 1
\end{pmatrix}"

This matrix is upper-triangular, and the diagonal entries are all non-zero. This implies the matrix is non-singular, and so the columns are linearly independent.

Thus, the set "\{1, (x+1), (x+1)^2\}" is a basis of "P_2."


2.

Using the technique of completing the square, we can factor the polynomial "f(x)=3+x+2x^2" as we like. Specifically,


"f(x)=3+x+2x^2"

"=2(x^2+2x+1)-4x-2+3+x"

"=2(x+1)^2-3(x+1)+4(1)"

Hence, we have the linear combination


"f(x)=3+x+2x^2=4(1)-3(x+1)+2(x+1)^2."